OpenStudy (anonymous):

y = 5x^2 conic sections parabola. have to find focus, directrix, and diameter. Answers are F(0,1/20) for focus, -1/20 for directrix, and 1/5 for diameter. please explain how they got these answers :)

OpenStudy (anonymous):

I know the setup is \[x ^{2}=20y\] but how are they getting the 1/4p times 5 to get that focus?

OpenStudy (jdoe0001):

\(\bf y=5x^2\implies (x-h)^2=4p(y-k)\qquad so\\ \quad \\ y=5x^2\implies (y-0)=5(x-0)^2\) (h, k) = vertex of the parabola p = distance from the vertex to the focus, or directrix, since both are equidistant to the vertex

OpenStudy (anonymous):

i see but the formulas my book gives me are x^2=4py and y^2=4px

OpenStudy (jdoe0001):

well... yes, those equations are correct for a parabola whose vertex is the origin

OpenStudy (anonymous):

heres the example they give that says should help solve this problem but to me its too simple to see where they they got those values and not enough written equations.

OpenStudy (anonymous):

look at example 4

OpenStudy (jdoe0001):

ok... anything wrong with that?

OpenStudy (anonymous):

it makes since but this problem im doing its alittle different so im not sure how they got the 1/20 for p

OpenStudy (jdoe0001):

1/2 you mean

OpenStudy (anonymous):

see heres the books answer to this problem. Number 15

OpenStudy (anonymous):

i know when you switch the formula from the original \[y = 5x^{2}\] you put it in the form \[x^{2} = 4py\] so that means it would be \[x^{2} = 20y\]

OpenStudy (jdoe0001):

\(\bf y=5x^2\implies (x-h)^2=4p(y-k)\qquad so\\ \quad \\ y=5x^2\implies (y-0)=5(x-0)^2\implies \cfrac{1}{5}(y-0)=(x-0)^2\\ \quad \\\implies (x-0)^2=\cfrac{1}{5}(y-0)\\ \quad \\ \begin{array}{llll} (x-h)^2=\color{red}{4p}(y-k)\\ (x-0)^2=\color{red}{\cfrac{1}{5}}(y-0) \end{array}\qquad\implies 4p= \cfrac{1}{5} \implies p=?\)

OpenStudy (anonymous):

so is my logic wrong?

OpenStudy (anonymous):

sadly we didnt get taught that formula for the vertex in this Trignomentry college course

OpenStudy (jdoe0001):

hmmm seems so.... once you put it in the form like above.... you can see what 4P equates to

OpenStudy (anonymous):

so to get that 1/5 did you just divide it out from the \[x^{2}\]

OpenStudy (jdoe0001):

you'd isolate the squared term, yes

OpenStudy (jdoe0001):

\(\bf x^2=4py\implies (x-0)^2=4p(y-0)\implies (x-h)^2=4p(y-k)\\ \quad \\ y^2=4px\implies (y-0)^2=4p(x-0)\implies (y-k)^2=4p(x-h)\) these two above, presume the vertex is at ( 0, 0) but the general forms are the ones with the (h, k) variables

OpenStudy (anonymous):

ok, so how do you solve for the diameter to get that 1/5 or is it just assumed?

OpenStudy (jdoe0001):

hmmm not sure on that one... since parabolas do not have diameters, circles do

OpenStudy (anonymous):

yea i know they call this a focal diameter which makes up the latus rectum

OpenStudy (anonymous):

thats why im confused :(

OpenStudy (anonymous):

there was an example on yahoo answers that got 1/4p thats why i was lost :( how they got that

OpenStudy (jdoe0001):

hmmm focal width... I see... focal width is 4P so 4 times P, \(\bf p = \cfrac{1}{20}\qquad \textit{focal width}\quad 4\cdot \cfrac{1}{20}\implies \cfrac{1}{5}\)

OpenStudy (anonymous):

i got another problem like this after this one so what do you recommend on approaching it because it is very similar

OpenStudy (jdoe0001):

same I'd think, change it to the general form first

OpenStudy (anonymous):

ok, so this one says x=-8y^2

OpenStudy (jdoe0001):

well, isolate the squared term first

OpenStudy (anonymous):

i made it -1/8x because variables by them selves divided are essentially 1 correct

OpenStudy (jdoe0001):

then \(\bf 4p=-\cfrac{1}{8}\)

OpenStudy (anonymous):

\[y^{2} = -1/8 x\]

OpenStudy (anonymous):

1/-32

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay i got them both right :) thank you!

OpenStudy (jdoe0001):

yw